OpenStudy (anonymous):

I need someone good at math Please tell me how to factor to solve: (2x-3)(3x+2)≤ 0. The answer to this equation is (-2/3) ≤ x ≤ (3/2)

OpenStudy (anonymous):

\(3x+2 \leq 0\) gives \(x \leq -2/3\)

OpenStudy (anonymous):

draw a table - 3 cols one for x<=-2/3, (-2/3) ≤ x ≤ (3/2) and x>=3/2 and a row each for 2x-3 3x+2 (2x-3)(3x+2) and find the sign after plugging in appropriate value you'll find that middle column will be neagative for bottom row